OpenStudy (anonymous):

Is it wrong if I say x^3-4 is the inverse of x^3+4?

OpenStudy (amistre64):

is it:\[x^{a+b}\]???

OpenStudy (anonymous):

no the -4 isn't part of the exponent.

OpenStudy (amistre64):

to bad; cause if it was they wouldve been inverses :)

OpenStudy (amistre64):

a good method for determining an inverse is to swap x and y parts; then solve for y y = x^3+4 ----------- x = y^3+4 x-4 = y^3 cbrt(x-4) = y

OpenStudy (anonymous):

what is cbrt?

OpenStudy (amistre64):

cube root

OpenStudy (amistre64):

\[\sqrt[3]{yada}\]

OpenStudy (anonymous):

OH :) I get it now! Thank you so much I really appreciate it I was having trouble with inverse!